# From Imitation to Intuition: Intrinsic Reasoning for Open-Instance Video Classification

### 11 Mar 2026 v1

@@ 4.4 Ablation Study & Analysis
We further analyze category-wise performance to understand where the proposed framework is most beneficial. As shown in Figure 6, our method consistently improves over the baseline across all categories, with an average gain of +4.50%. The largest gains are observed on Gambling (+8.33%), Alcohol (+6.19%), Personal Information (+5.60%), and Tobacco & Nicotine (+5.42%), indicating that intrinsic reasoning is particularly helpful in categories that require more nuanced contextual understanding and intent-sensitive judgment. Moderate gains are also achieved on High Risk Driving (+4.43%) and Designated Political Entities (+3.29%), suggesting improved robustness in semantically complex or safety-sensitive cases. By contrast, the gains are smaller on Fraud and Scams (+0.53%) and Adult Sexualized Behaviors (+1.88%), where the baseline is already relatively strong or the decision depends more directly on explicit visual evidence. Overall, these results show that the benefits of intrinsic reasoning are consistent across categories, with the strongest improvements appearing in cases that involve subtle semantic distinctions.

### 13 Mar 2026 v2

@@ 4.4 Ablation Study & Analysis
We further analyze category-wise performance to understand where the proposed framework is most beneficial. As shown in Figure 6, our method consistently improves over the baseline across most categories. The largest gains appear on label #6 and label #7, where our approach significantly outperforms the Qwen-SFT baseline, indicating that intrinsic reasoning is particularly beneficial for categories requiring more contextual interpretation. Noticeable improvements are also observed on label #1 and label #2, where the performance gap remains consistently positive. Moderate gains are achieved on label #3, label #8, and label #9, suggesting improved robustness in more ambiguous scenarios. By contrast, the improvement is relatively smaller on label #4 and label #5, where the baseline already performs competitively and the decision boundaries may rely more on explicit visual evidence. Overall, the results demonstrate that intrinsic reasoning provides consistent benefits across diverse categories, with the most substantial improvements appearing in cases that require deeper semantic understanding.